Home‎ > ‎Reports‎ > ‎Crew110a‎ > ‎

day05


Crew #110A Log Book for January 4, 2012

Photos of the day





Commander's Check-In Report

Date and time: 4 January 2012, 18:00

Written by: Julie Mason, Crew 110, Commander

Crew Physical Status: Great

Time Departed/Returned from EVA: EVA 1 10:15-11:45, EVA 2 13:15-15:00;

Telescope EVA (Last Night): 19:00 – 01:45

Brief Narrative of Field Mission Results:

Last night the crew took 3 minute videos of Jupiter at the Musk Mars Desert Observatory at one hour increments. The crew processed images using RegiStax 6. This data will go towards our atmospheric study. On off hours we took videos of the moon and observed star clusters like the Pleiads, better known as the Seven Sisters.

The crew observed the Quadrantid Meteor Shower with the naked eye from 1:00am- 3:30 in the Northeast sky.

Crew members continued to exercise in the habitat while not on EVA. Today consisted of a little cardio and an abdominal work-out.

During EVA 1, two crew members created outreach videos. One showed the process of putting on a space suit at MDRS. The other explored the terrain around the habitat. More videos will be made indoors while not on EVA.

During EVA 2, three crew members worked on the 3D imaging of the habitat and surround terrain.

Engineering/Hab Maintenance: Spark on gas stove needs to be fixed

Report Transmission Schedule:

Plans for Tomorrow: EVA 1—In the morning, the crew plans to make more outreach videos. EVA 2— 3D Imaging Project, EVA 3—Observatory, Atmospheric Study

Inventory: Good

Support Requested: None

Miscellaneous: The crew found last year’s CD: A Mars New Year’s Eve remix. It was a good CD to play during our workout today.

Commander's Report

Date and time: 4 January 2012 17:18

Written by Julie Mason, Crew 110A, Commander

Lyndsey and I worked on outreach videos this morning. We made one talking about everyday life upstairs in the habitat, another showing the details of suiting up for an EVA, and another exiting the habitat and exploring the terrain. These will be edited and audio will be added to the exploration videos.

This afternoon, Aaron, Mark, and I finished imaging the last quarter of the habitat. The crew is anxiously awaiting the finalized 3D model that Aaron is completing.

Today was the first day four people exercised in the habitat at the same time. There is plenty of room in the habitat to accomplish this. It’s quite spacious.

Tonight, Lyndsey, Sam, Will, and I will take shifts imaging Sirius at the observatory in increments over the course of the night.

3D Model Project Report

Date and time: 4 January 2012, 18:45

Written by Aaron Olson, Mission Specialist

During the fifth day of Crew 110A’s stay at the MDRS site, Mark, Julie and I completed taking images around the Hab. I will continue processing the images and will start scaling of the full model. Based on the results of the processed model, we will take additional photos to get the best model possible.

EVA 4 Report

Date and time: 4 January 2012, 23:30

Written By: Lyndsey Bankers, Crew 110a, Mission Specialist


Time of EVA: 10:15-11:45

Crew Members: Julie Mason and Lyndsey Bankers

Site Location: Near the Hab

Transit Mode: ATV’s


Objectives: To create outreach videos


Results: We recorded a few videos that we plan on editing to put

different pieces of the videos together.

Lessons Learned: It’s hard to make videos when you are wearing the

space suits and can’t talk.


EVA 5 Report

Date and time: 4 January 23:30

Written by: Julie Mason,Crew 110A, Commander

Time: 13:15-15:00

Crew Members: Aaron Olson, Mark Ruff, Julie Mason

Site Location: Within 250 ft of habitat

Transit Mode: Foot

Objectives: Measure distances for 3D images, continue imaging Hab and surrounding terrain

Results: Collected the last quarter of the images needed for main 3D model. Model will be worked on tonight in the Hab, and if more photos are needed, they will be taken tomorrow.

Lessons Learned: We switched to marking points with GPS. It’s more efficient.

EVA 6 Report

Date and time: 4 January 23:30

Written by: Julie Mason, Crew 110A, Commander

Time: 19:00-01:45

Crew Members: Will Yu, Lyndsey Bankers, Sam Marron, Julie Mason

Site Location: Musk Mars Desert Obervatory

Transit Mode: Foot

Objectives: The purpose of this study is to observe a set of stars or a planet at different positions in the sky. Assuming the atmosphere is constant for all azimuths, we will do a linear regression to determine the amount of extinguishing material in the atmosphere. On Mars, this could be used to monitor the amount of aerosols in the atmosphere at night.

Results: Jupiter was imaged overnight. Frames will be compiled in Registax during the day tomorrow.

Lessons Learned: As Jupiter moved across the sky, the need to refocus the telescope became more apparent each hour. We will image stars tomorrow night.

Engineering Report

Date and time: 04 January, 2012, 18:30 MST
Written by Will Yu, Crew 110A, Chief Engineer.

Generator/Electricity:
Did you have any electricity outage? (Yes/No): NO

Fuel Status:
Diesel (Full/0.75/0.5/0.25/Empty): .75
Propane Tank (Full/0.75/0.5/0.25/Empty): .75
Gasoline Tank (Full/0.75/0.5/0.25/Empty): .50

Oil:
ATV Oil Quantity (# of Quarts in storage): 0
Car Oil Quantity (# of Quarts in storage): N/A

Water Status:
Outside Potable Water Tank Level (inches from bottom): 17
Trailer Potable Water Tank Level (inches from bottom): 26
Hab Potable Water Tank Level (inches from bottom): 11
Potable Water Meter Reading: 48591.7
Into GreenHab Meter Reading: N/A
Notes/Comments:

GreenHab:
Condition of plants in Tank 1: N/A
Condition of plants in Tank 2: N/A
Condition of plants in Tank 3: N/A

Observatory:
Telescope Used? (Yes/No): YES
Observatory Notes/Comments: N/A

Transportation:
Hab Car (NOT RENTAL)
Used (yes/no): NO
Oil Checked (yes/no): NO
Tire Status: Good
Odometer at end of day: 125488.9
Notes/Comments on Hab Car:

ATV 1 (Opportunity):
Used: YES
Oil Checked: NO
Fuel Consumed: 1 gal
Tire Status: Good - 4psi
Today's Run Time: 01:00
Season's Run Time: 45:50
Comments:

ATV 2 (Spirit):
Used: NO
Oil Checked: NO
Fuel Consumed: 0 gal
Tire Status: Good - 4psi
Today's Run Time: 00:00
Season's Run Time: 37:53
Comments:

ATV 3 (Viking I):
Used: YES
Oil Checked: NO
Fuel Consumed: 1 gal
Tire Status: Good
Today's Run Time: 01:00
Season's Run Time: 42:35
Comments:

Heating and Ventilation:
Thermostat setting upstairs (Farenheit): Upstairs furnace not turned on since new downstairs furnace was enabled
Actual temperature upstairs (Farenheit): 81 F

Thermostat setting downstairs (Farenheit): Downstairs furnace was left on overnight (00:00 – 13:00).
Actual temperature downstairs (Farenheit): 72 F

Computers/Networking Infrastructure:
Notes/comments on internet/computers?:

General Engineering Notes/Comments:
~ Furnace was left on for too long, downstairs hab reached 72 degrees
~ Helmet 5 had an existing crack which has no propagated a little further
~ Spark on the stove no longer lights the burners. Manual lighting of the stove required
~ Need ATV oil for available for storage

Journalist Report

Date and time: 4 January 2012 7:00 PM

Written by: Mark Ruffh

Today Lyndsey and Julie woke up early to go out on an EVA to record some footage for outreach materials. Sam, Will, Aaron and started getting up as they prepared to leave. Once they made it back to the Hab, we all prepared a short lunch, before Aaron, Julie and I went out to take the last round of photographs for the 3D imaging study. After that, Aaron and I took the ATV’s out to survey some good spots for some later photography. After an excellent dinner of soup and tamales, the crew has split up to work on reports, and work with the observatory. Not much else to report. Hopefully there will be some good photos with the reports tomorrow.

Chef's Report


Date and Time: 4 January 2011

Written by: Will Yu, Crew 110A


1a) Today was a non-cooking day (pre-prepared)

1b) Was today a special day celebrated at the main meal with special food or activity? If so, what was special about today and what food(s) or activities were prepared to mark the celebration?
No

2) Backpacker's Pantry Western Tamale Pie, chicken noodle soup, bacon, milk

3) 3 bowls of tamale pie and 3 bowls of soup leftover stored in fridge

4) All food prepared were simply rehydrated, aside from the bacon, which was microwaved

5) No lessons learned

6) No additional comments or questions